OpenStudy (anonymous):

A sequence has its first term equal to 3, and each term of the sequence is obtained by adding 5 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 3 and f(n) = f(n - 1) + 5; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3; n > 1 f(1) = 3 and f(n) = f(n - 1) + 5n; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3n; n > 1

OpenStudy (anonymous):

Since the first term is 3, you can eliminate the choices with \(f(1)=5\).

OpenStudy (anonymous):

Beyond that, you can choose between the others by plugging in for \(n\). According to the definition of this recursion, the next term will add 5 to the previous term. So if the first term is \(f(1)=3\), then the second term must be \(f(2)=3+5=8\), or \(f(2)=f(1)+5\). What does that tell you?

OpenStudy (anonymous):

that for every F(x), its 3+5+whatever the last number is

OpenStudy (anonymous):

Not quite, that would mean the sequence is defined by \[\begin{cases}f(1)=3\\f(n)=f(n-1)+8\end{cases}\] which is not one of the options.

OpenStudy (anonymous):

ohh, so just plus 5 every time then?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@nikosis01 can you help real quick?

OpenStudy (nikosis01):

sure

OpenStudy (anonymous):

@KryoWolf yes just plus 5

OpenStudy (anonymous):

ok thanks sith

OpenStudy (anonymous):

so if f(2)=3+5=8, then f(3)=8+5=13

OpenStudy (anonymous):

so C @SithsAndGiggles

OpenStudy (anonymous):

No, not C. C says \[f(n)=f(n-1)+5n\] But that would mean \[f(2)=f(2-1)+5(2)=f(1)+10=13\not=8\]

OpenStudy (anonymous):

ok, so its A

OpenStudy (anonymous):

yes